OpenStudy (r.hamza17):

The pth term of an Arithmetic progression is 1/q and qth term is 1/p. The sum of the pqth term is: a) 1/2(pq+1) b) 1/2(pq-1) c) pq+1 d) pq-1

OpenStudy (trojanpoem):

\[a_{1} + d(p - 1) = \frac{ 1 }{ q }\] \[a_{1} + d(q-1) = \frac{ 1 }{ p }\]

OpenStudy (r.hamza17):

We dont need to but the sum of pqth term we just need to check what is the sum of the pqth term.

OpenStudy (trojanpoem):

-a1 - d(p-1) = -1/q a1 + d(q-1) = 1/p -d(p-1) + d(q-1) = -1/q + 1/p -dp +d +qd -d = -1/q + 1/p d(p-q) = p - q /pq d = 1/pq -> common difference a1 = 1/q - 1/pq * p + 1/pq a1 = 1/q - 1/q +1/pq a1 = 1/pq

OpenStudy (trojanpoem):

\[a_{n} = \frac{ 1 }{ pq } + \frac{ 1 }{ pq } (n-1) = \frac{ n }{ pq }\]

OpenStudy (trojanpoem):

Now we find the sum.

OpenStudy (trojanpoem):

\[a_{n} = \frac{ n }{ 2 } ( 2 a + d(n-1))\]

OpenStudy (trojanpoem):

\[a_{pq} = \frac{ pq }{ 2 } ( \frac{ 2 }{ pq } - \frac{ pq - 1 }{ pq })\]

OpenStudy (trojanpoem):

\[\frac{ pq }{ 2 }(\frac{ 2 + pq - 1 }{ pq }) = 1 + pg/2 = 0.5(pg+1)\]

OpenStudy (trojanpoem):

Correct the negative sign in the step before this.

OpenStudy (r.hamza17):

Thanx alot mate

OpenStudy (trojanpoem):

Any time.